How do we model a falling object coming to rest when hitting another object using a differential equation? eg a ball coming to rest after hitting a mattress.
Answer #1

3 F=-ng-kv md=-ng-kv vlo)=0,ko dt Fy=mg dt W dt = -t . gtk v c let u= gt kv M du = k dun gt tooktim g+ kV – ge 312 Horn = g( cut/-1) - my (ext|m-1) V (t) = – mg (1-ekting Terminal velocity